A new community center is being built in Safe Harbor. The perimeter of the rectangular playing field is 444 yards. The length of the field is 2 yards less than triple the width. What are the dimensions of the playing field?
Answer by JulietG(1812)

Remember that a perimeter is like a fence around. To fence an area, you need twice the width and twice the length.
P = 2W + 2L
What do we know from the question?
P = 444
L = 3W - 2
Plug those values into the first equation.
444 = 2W + 2(3W-2)
444 = 2W + 6W - 4
448 = 8W
Divide
56 = W
L = 3W - 2
L = 3(56) - 2
L = 168-2
L = 166
Let make sure that's correct by putting the values back into the original equation.
P = 2W + 2L
444 = 2(56) + 2(166)
444 = 112 + 332
444 = 444
Success!